OpenStudy (anonymous):

x + 9 = 5(4x – 2) (2 points) elevennineteenths –1 1 negativeonenineteenth

OpenStudy (anonymous):

x+9=5(4x-2) x+9=20x-10 9+10=20x-x 19=19x 19/19=x 1=x Option 3rd.

OpenStudy (anonymous):

x+9=20x-10 9+10=20x-x 19=19x x=1

OpenStudy (anonymous):

okay thanks:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Not a problem :)